OpenStudy (anonymous):

Is the line through points P(–7, 2) and Q(7, –6) parallel to the line through points R(–1, –5) and S(–3, –1)? Explain. a. No; one line has zero slope, the other has no slope. b. Yes; the lines are both vertical. c. Yes; the lines have equal slopes. d. No; the lines have unequal slopes.

OpenStudy (anonymous):

your job is to compute two slopes do you know how to do that?

OpenStudy (anonymous):

I tried but I cant figure out which number goes where. How do you do that?

OpenStudy (anonymous):

\[\frac{y_2-y_1}{x_2-x_1}\] with say \(x_1=-7,y_1=2,x_2=7,y_2=-6\)

OpenStudy (anonymous):

you get as a first step \[\frac{-6-2}{7-(-7)}\] then compute via \[\frac{-8}{14}=-\frac{4}{7}\]

OpenStudy (anonymous):

then repeat for the second pair if you get the same number, they are parallel

OpenStudy (anonymous):

neither line is vertical , because the \(x\) coordinates are different for both of them, so forget that answer

OpenStudy (anonymous):

So they ARE parallel, right?

OpenStudy (anonymous):

i don't know i didn't compute the second one let me check

OpenStudy (anonymous):

\[\frac{-1-(-5)}{-3-(-1)}\] no i don't think you get the same answer

OpenStudy (anonymous):

i get \[\frac{4}{-2}=-2\]for the second one

OpenStudy (anonymous):

Dang it, I did it wrong. Hold on. So its no because they have unequal slopes?

OpenStudy (anonymous):

yes the slopes are not equal, so not parallel

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw